Inactive Tutor answered 10/11/20
We know he bought an ice cream on Monday
So the possibilities are
Tuesday Wednesday Tues prob Wed prob Prob of both
Ice Drink 0.70 0.30 0.7 x 0.3 = 0.21
Ice Ice 0.70 0.70 0.49
Drink Ice 0.30 0.40 0.12
Drink Drink 0.30 0.60 0.18
So probability (ice on Wed) = 0.49 + 0.12 = 0.61
Can also use a probability tree diagram here
Michael C.
How would the Conditional Probability Formula be used for this situation?10/13/21
Daniel R.
Thank you 👍😊10/11/20